# Sparse Autoencoders Reveal Universal Feature Spaces Across Large Language Models

### v3

@@ Acknowledgments
We thank the Torr Vision Group at the University of Oxford for providing compute and hosting Michael Lan during the internship. We thank Clement Neo, Luke Marks, Minseon Kim, Lovis Heindrich, Tingchen Fu, Constantin Venhoff, Hazel Kim, Will Lin, Alasdair Paren, Christian Schroeder de Witt, Kiho Park, and Max Klabunde for providing helpful feedback. Clement Neo helped with the design of Figure 1.
@@ Appendix B Trained SAE Metric Evaluations
In general, all of our custom trained SAEs on Pythia-70m and on Pythia-160m have similar metrics as the ones reported for the SAEs in this section. We evaluate the SAEs on SAEBench (Karvonen et al., 2024).
For the random SAEs trained on identical conditions as the non-random SAEs mentioned above, but on a randomized Pythia-70m with the embedding layer not being randomized, we obtain explained variances of between 0.63 to 0.80 ($\mu=0.71,\sigma=0.006$), cosine similarities  1, cross entropy losses of 1, and L1 norm between 27.38 to 38.25. These results are similar to the results for Pythia-70m-deduped SAEs found by Heap et al. (2025), and indicates SAEs with good reconstruction and sparsity metrics.
@@
Appendix C Results from More Models
C.1 Gemma-2-2B vs Gemma-2-9B Results
We compare Gemma-2-2B vs Gemma-2-9B at several layers of each model, which each have 25 and 41 layers, respectively. Each SAE has a width of 16384 features and uses JumpReLU (Rajamanoharan et al., 2024b), and is hosted on SAELens. We obtain activations from OpenWebText using 150 samples with a max sequence length of 150, for a total of 22.5k tokens. Figure 4 shows SVCCA and RSA scores after filtering keywords and for 1-1. On average, around 20% of feature pairs are kept after filtering. In particular, the L11 vs L21 comparison achieves an SVCCA score of 0.70 (while mean random pairing has a score of 0.009) and an RSA score of 0.195 (while mean random pairing has a score of $4.38\times 10^{-4}$).
C.2 Base vs Fine Tuned Results
We compare Gemma-2-2B vs Gemma-2-9B-Instruct, a fine tuned version, at Layers 9, 20, and 31 of each model. Each SAE has a width of 16384 features. Only the pretrained SAEs for layers 9, 20 and 31 were publically available online and hosted on SAELens. We obtain activations from the RedPajamas dataset (Weber et al., 2024) using 150 samples with a max sequence length of 150, for a total of 22.5k tokens. On average, 22.2% of feature pairs are kept after filtering, with the range between 16.4% and 33.5%. Figure 5 shows that closer layers had higher SVCCA scores, while RSA scores were low except for matching layers. This is consistent with previous work which found that SAE features can transfer across base to fine-tuned language models (Kissane et al., 2024a; Kutsyk et al., 2024).
@@ C.2 Base vs Fine Tuned Results
C.3 Llama 3 Results
We compare Llama 3-8B-Instruct and Llama 3.1-8B at each model’s Layer 25 residual stream, a later layer in both models. The Llama 3 SAE has a width of 65536 features and uses Gated ReLU (Rajamanoharan et al., 2024a), while the Llama 3.1 SAE has a width of 32768 features and uses JumpReLU (Rajamanoharan et al., 2024b). We obtain activations using 100 samples with a max sequence length of 100, for a total of 10000 tokens. After filtering, this comparison achieves an SVCCA score of 0.3, comparable to the score of later layers in both Gemma and Pythia.
@@ Appendix D Baseline Comparisons to SAEs Trained on LLMs with Randomized Weights
Recent research has demonstrated that training SAEs on randomly initialized transformers—where parameters are independently sampled from a Gaussian distribution rather than learned from text data—yields latent representations that are similarly interpretable to those from trained transformers (Heap et al., 2025). One caveat is that the interpretable features of SAEs trained on "randomized" LLMs tend to activate on simple features, such as ones that activate on specific single tokens instead of more general concepts, rather than complex abstract features learned by middle to later layer SAEs. The authors propose that there may be cases where "simple" interpretable latents may arise from the inherent sparsity of the text data used to train language models, rather than from uncovering the computational structure embedded within the underlying model. Thus, we test whether the feature similarities found across LLMs are due to the sparsity of the training data or due to the computational structure. We run experiments comparing the representational similarity of SAE features trained on randomized LLMs to SAE trained on the original LLMs. We call the former "random SAEs", and the latter "non-random SAEs".
We find that, relative to other across model experiments, there is little similarity between the non-random SAEs and random SAEs, which suggests that our method is demonstrating SAE feature space similarity that is capturing the model’s computational structure, and is not just capturing similarity due to interpretability from the inherent sparity of the text data. In particular, not even the same layers have medium similarity. However, the small amounts of similarity do suggest that SAEs trained on even randomized LLMs can capture features in the data, which is consistent with the interpretable "simple, low information" features, such as those that just activate on single tokens, that were also present in random SAEs found by Heap et al. (2025).
SAE Models. For Pythia-70m, we compare residual stream layer SAEs trained a randomized Pythia-70m, to SAEs trained on the original Pythia-70m. Both SAEs were trained using 100 million tokens from the RedPajama dataset (Weber et al., 2024), with an expansion factor of 64 and a Top-K of 32. The embedding layer was not randomized. We evaluate the SAEs on SAEBench (Karvonen et al., 2024) and report their results in Section §B. Overall, all SAEs obtained good reconstruction and sparsity metric scores.
@@ E.1 Comparing the same SAE to themselves
This is consistent with recent results which found that SAEs do not learn consistent features (Leask et al., 2025; Hindupur et al., 2025). Thus, while we show that certain subspaces learned by LLMs have high similarity, as revealed by SAEs, these subspaces may not always be consistently captured by arbitrary SAEs. This demonstrates a limitation by SAEs- they can capture different features that are shared across models, but cannot do this consistently. Thus, future work can attempt to address this limitation to develop SAEs that capture features more consistently. Indeed, recent work has already developed "universal SAEs" that can better capture features found across models (Thasarathan et al., 2025).
@@ E.2 Comparing SAEs trained with different seeds
As mentioned before in Section §E.1, this is consistent with recent results which found that SAEs do not learn consistent features (Leask et al., 2025; Hindupur et al., 2025). Despite not finding the same features, which can be viewed as a form of strong universality, these SAEs still show notable similarity. This can be viewed as a form of weak universality; see Appendix H for further discussion on this topic.
@@
SAE Training Parameters. We train two SAEs at different seeds for a TinyStories-1L-21M model (Eldan, 2023). Both SAEs are trained at 30k steps with expansion factor of 16, for a 16384 total features, on a TinyStories dataset, and using ReLU as the activation function. We use a batch size of 4096 tokens and train with an Adam optimizer, applying a constant learning rate of $1\times 10^{-5}$. The learning rate follows a warm-up-free schedule with a decay period spanning 20% of the total training steps. L1 sparsity regularization is applied with a coefficient of 5, and its effect gradually increases over the first 5% of training steps. We obtain SAE activations using 500 samples from the TinyStories dataset with a max sequence length of 128 (for a total of 64000 tokens).
@@ E.3 Comparing SAEs with different widths
In general, we do not find noticeable trends in scores as we vary SAE widths. This may be because each SAE varies in the features they capture Leask et al. (2025), so while they capture feature similarities across models, the similar features that they do capture are not the same for each SAE. We did not evaluate for feature splitting or absorption (Chanin et al., 2024).
@@ E.4 Comparing SAEs trained on different datasets
Figure 12 shows that the two sets of SAEs show some similarities, such as at Layers 1, 3, and 5, but also notable dissimilarities, such as having nearly no similarities at Layers 2 and 4. These discrepencies may be due to the difference in training data, as it has been shown that SAEs are highly dataset dependent (Kissane et al., 2024b). However, despite these discrepencies in individual features, the feature subspaces we compare appears to be highly similar, which may suggest that while the specific values captured by feature weight vectors may differ, these values lie in similar latent space regions that capture similar semantics.
@@ Appendix H Universal Vs True Features
Our definition of universal features is based on Claim 3 of "Zoom In: An Introduction to Circuits", which defines universality as studying the extent to which analogous features and circuits form across models and tasks (Olah et al., 2020). Previous work has investigated similar claims by showing evidence for highly correlated neurons (Yosinski et al., 2014; Gurnee et al., 2024) and similar feature space representations at hidden layers (Kornblith et al., 2019; Klabunde et al., 2023). Using these definitions of universality, our results find both highly correlated features and similar feature representations in SAEs.
@@
On the other hand, “true features” (Bricken et al., 2023) can be interpreted as atomic linear directions, such that activations are a linear combination of "true feature" directions (Till, 2023), though this term has not been formally defined yet. This definition for features imposes stricter requirements which do not just rely on high correlation and similar representations, but define "true features" as linear directions that model the same concepts, assuming they exist in ground truth. Previous work has shown that SAEs may capture ground truth features Sharkey et al. (2022), while recent research has provided evidence that SAE features are not atomic (Bussmann et al., 2024), which suggests SAEs may not be sufficient to capture "true features" if they exist.
@@
Chughtai et al. (2023a) define universality in terms of weak and strong universality. Strong universality claims that models trained in similar ways will develop the same features and circuits, whereas weak universality argues that fundamental principles, such as similar functionalities, are learned across models, but their exact implementations may vary considerably. From this perspective, strong universality would correspond more with "true features", while our paper would align closer to measuring weak universality in terms of similar feature space correlations. However, we also define analogous feature universality as a property between strong and weak universality which hypothesizes that while overall representations may greatly differ, there may be transformations that can be applied to a subset of these representations to better reveal their similarity.
@@
The Importance of Analogous Features. While features across models may not capture the same ground truth feature representations, having analogous feature spaces could suggest that there may be transformations or mappings between the representation spaces of models which reveal their similarity. As such, this has implications for transfer learning and model stitching in which representations, including cross-layer circuits, from one model may be transferred to another (possibly larger) model through a mapping model. This may allow for functionality transfer. These studies may also have implications for furthering studies in neuroscience, as past research has been conducted on measuring the extent of alignment across biological and artificial neural networks (Goldstein et al., 2024), and previous work has also found high-low feature detectors in both vision models (Schubert et al., 2021) and mouse brains (Ding et al., 2023).
@@
This paper seeks to measure the extent of universality in terms of "analogous features in representational space", and unlike Huh et al. (2024), does not claim that models converge towards finding universal features. This follows previous papers which also measured the extent to which different neural networks learn the same representation (Wang et al., 2018a). We provide evidence that shows there are non-negligible signals of universality captured by subsets of SAE feature spaces across models, but also claim that many feature space subsets in SAEs are dissimilar in terms of their representations and relations in feature space. Still, since this may be due to the limitations of SAEs in capturing features (Engels et al., 2024b), these results do not conclude that in these dissimilar areas, models do not share universal features, and further work must be done in this area to draw stronger conclusions.
@@ Appendix I Illustrated Steps of Comparison Method
Figure 19 demonstrates steps 1 to 3 to carry out our similarity comparison experiments.
@@
Appendix L Concept Categories List and Further Analysis
The keywords used for each concept group for the semantic subspace experiments are given in Table 4. We generate these keywords using GPT-4 Bubeck et al. (2023) using the prompt "give a list of N single token keywords as a python list that are related to the concept of C", such that N is a number (we use N=50 or 100) and C is a concept like Emotions. We then manually filter these lists to avoid using keywords which have meanings outside of the concept group (eg. "even" does not largely mean "divisible by two" because it is often used to mean "even if…"). We also aim to select keywords which are single-tokens with no spaces, and are case-insensitive.  [^4] However, not all the tokens in the Table 4 are single token.
When searching for keyword matches from each feature’s list of top 5 highest activating tokens, we use keyword matches that avoid features with dataset samples which use the keywords in compound words. For instance, when searching for features that activate on the token "king" from the People/Roles concept, we avoid using features that activate on "king" when "king" is a part of the word "seeking", as that is not related to the People/Roles concept.
We perform interpretability experiments to check which keywords of the features kept are activated on. We find that for many feature pairs, they activate on the same keyword and are monosemantic. Most keywords in each concept group are not kept. Additionally, for the layer pairs we checked on, after filtering there were only a few keywords that multiplied features fired on. This shows that the high similarity is not because the same keyword is over-represented in a space.
For instance, for Layer 3 in Pythia-70m vs Layer 5 in Pythia-160m, and for the concept category of Emotions, we find which keywords from the category appear in the top 5 activating tokens of each of the Pythia-70m features from the 24 feature mappings described in Table 1. We only count a feature if it appears once in the top 5 of a feature’s top activating tokens; if three out of five of a feature’s top activating tokens are "smile", then smile is only counted once. The results for Models A and B are given in Table 3. Not only are their counts similar, indicating similar features, but there is not a great over-representation of features. There are a total of 24 features, and a total of 25 keywords (as some features can activate on multiple keywords in their top 5). We note that even if a feature fires for a keyword does not mean that feature’s purpose is only to "recognize that keyword", as dataset examples as just one indicator of what a feature’s purpose is (which still can have multiple, hard-to-interpret purposes).

### v4

@@ Acknowledgments
We thank the Torr Vision Group at the University of Oxford for providing compute and hosting Michael Lan during the internship. We thank Clement Neo, Luke Marks, Minseon Kim, Lovis Heindrich, Tingchen Fu, Constantin Venhoff, Hazel Kim, Will Lin, Alasdair Paren, Christian Schroeder de Witt, Kiho Park, and Max Klabunde for providing helpful feedback. Clement Neo helped design Figure 1.
@@ Appendix B Trained SAE Metric Evaluations
In general, all of our custom trained SAEs on Pythia-70m and on Pythia-160m have similar metrics as the ones reported for the SAEs in this section. We evaluate the SAEs on SAEBench [Karvonen et al., 2024].
For the random SAEs trained on identical conditions as the non-random SAEs mentioned above, but on a randomized Pythia-70m with the embedding layer not being randomized, we obtain explained variances of between 0.63 to 0.80 ($\mu=0.71,\sigma=0.006$), cosine similarities  1, cross entropy losses of 1, and L1 norm between 27.38 to 38.25. These results are similar to the results for Pythia-70m-deduped SAEs found by Heap et al. [2025], and indicates SAEs with good reconstruction and sparsity metrics.
@@
Appendix C Base vs Fine Tuned Results
We compare Gemma-2-2B vs Gemma-2-9B-Instruct, a fine tuned version, at Layers 9, 20, and 31 of each model. Each SAE has a width of 16384 features. Only the pretrained SAEs for layers 9, 20 and 31 were publically available online and hosted on SAELens. We obtain activations from the RedPajamas dataset [Weber et al., 2024] using 150 samples with a max sequence length of 150, for a total of 22.5k tokens. On average, 22.2% of feature pairs are kept after filtering, with the range between 16.4% and 33.5%. Figure 5 shows that closer layers had higher SVCCA scores, while RSA scores were low except for matching layers. This is consistent with previous work which found that SAE features can transfer across base to fine-tuned language models [Kissane et al., 2024a, Kutsyk et al., 2024].
@@ Appendix D Baseline Comparisons to SAEs Trained on LLMs with Randomized Weights
Recent research has demonstrated that training SAEs on randomly initialized transformers—where parameters are independently sampled from a Gaussian distribution rather than learned from text data—yields latent representations that are similarly interpretable to those from trained transformers [Heap et al., 2025]. One caveat is that the interpretable features of SAEs trained on "randomized" LLMs tend to activate on simple features, such as ones that activate on specific single tokens instead of more general concepts, rather than complex abstract features learned by middle to later layer SAEs. The authors propose that there may be cases where "simple" interpretable latents may arise from the inherent sparsity of the text data used to train language models, rather than from uncovering the computational structure embedded within the underlying model. Thus, we test whether the feature similarities found across LLMs are due to the sparsity of the training data or due to the computational structure. We run experiments comparing the representational similarity of SAE features trained on randomized LLMs to SAE trained on the original LLMs. We call the former "random SAEs", and the latter "non-random SAEs".
We find that, relative to other across model experiments, there is little similarity between the non-random SAEs and random SAEs, which suggests that our method is demonstrating SAE feature space similarity that is capturing the model’s computational structure, and is not just capturing similarity due to interpretability from the inherent sparity of the text data. In particular, not even the same layers have medium similarity. However, the small amounts of similarity do suggest that SAEs trained on even randomized LLMs can capture features in the data, which is consistent with the interpretable "simple, low information" features, such as those that just activate on single tokens, that were also present in random SAEs found by Heap et al. [2025].
SAE Models. For Pythia-70m, we compare residual stream layer SAEs trained on a randomized Pythia-70m, to SAEs trained on the original Pythia-70m. Both SAEs were trained using 100 million tokens from the RedPajama dataset [Weber et al., 2024], with an expansion factor of 64 and a Top-K of 32. The embedding layer was not randomized. We evaluate the SAEs on SAEBench [Karvonen et al., 2024] and report their results in Section §B. Overall, all SAEs obtained good reconstruction and sparsity metric scores.
@@ E.1 Comparing the same SAE to themselves
This is consistent with recent results which found that SAEs do not learn consistent features [Leask et al., 2025, Hindupur et al., 2025]. Thus, while we show that certain subspaces learned by LLMs have high similarity, as revealed by SAEs, these subspaces may not always be consistently captured by arbitrary SAEs. This demonstrates a limitation of SAEs- they can capture different features that are shared across models, but cannot do this consistently. Thus, future work can attempt to address this limitation to develop SAEs that capture features more consistently. Indeed, recent work has already developed "universal SAEs" that can better capture features found across models [Thasarathan et al., 2025].
@@ E.2 Comparing SAEs trained with different seeds
As mentioned before in Section §E.1, this is consistent with recent results which found that SAEs do not learn consistent features [Leask et al., 2025, Hindupur et al., 2025]. Despite not finding the same features, which can be viewed as a form of strong universality, these SAEs still show notable similarity. This can be viewed as a form of weak universality; see Appendix H for further discussion on this topic.
@@
SAE Training Parameters. We train two SAEs at different seeds for a TinyStories-1L-21M model [Eldan, 2023]. Both SAEs are trained at 30k steps with expansion factor of 16, for a 16384 total features, on a TinyStories dataset, and using ReLU as the activation function. We use a batch size of 4096 tokens and train with an Adam optimizer, applying a constant learning rate of $1\times 10^{-5}$. The learning rate follows a warm-up-free schedule with a decay period spanning 20% of the total training steps. L1 sparsity regularization is applied with a coefficient of 5, and its effect gradually increases over the first 5% of training steps. We obtain SAE activations using 500 samples from the TinyStories dataset with a max sequence length of 128 (for a total of 64000 tokens).
@@ E.3 Comparing SAEs with different widths
In general, we do not find noticeable trends in scores as we vary SAE widths. This may be because each SAE varies in the features they capture Leask et al. [2025], so while they capture feature similarities across models, the similar features that they do capture are not the same for each SAE. We did not evaluate for feature splitting or absorption [Chanin et al., 2024].
@@ E.4 Comparing SAEs trained on different datasets
Figure 12 shows that the two sets of SAEs show some similarities, such as at Layers 1, 3, and 5, but also notable dissimilarities, such as having nearly no similarities at Layers 2 and 4. These discrepencies may be due to the difference in training data, as it has been shown that SAEs are highly dataset dependent [Kissane et al., 2024b]. However, despite these discrepencies in individual features, the feature subspaces we compare appears to be highly similar, which may suggest that while the specific values captured by feature weight vectors may differ, these values lie in similar latent space regions that capture similar semantics.
@@ Appendix H Universal Vs True Features
Our definition of universal features is based on Claim 3 of "Zoom In: An Introduction to Circuits", which defines universality as studying the extent to which analogous features and circuits form across models and tasks [Olah et al., 2020]. Previous work has investigated similar claims by showing evidence for highly correlated neurons [Yosinski et al., 2014, Gurnee et al., 2024] and similar feature space representations at hidden layers [Kornblith et al., 2019, Klabunde et al., 2023]. Using these definitions of universality, our results find both highly correlated features and similar feature representations in SAEs.
@@
On the other hand, “true features” [Bricken et al., 2023] can be interpreted as atomic linear directions, such that activations are a linear combination of "true feature" directions [Till, 2023], though this definition is still subject to change in future works. This definition for features imposes stricter requirements which do not just rely on high correlation and similar representations, but define "true features" as linear directions that model the same concepts, assuming they exist in ground truth. Previous work has shown that SAEs may capture ground truth features Sharkey et al. [2022], and provided evidence that SAE features are not atomic [Leask et al., 2025], which suggests SAEs may not be sufficient to capture "true features" if they exist.
@@
Chughtai et al. [2023] define universality in terms of weak and strong universality. Strong universality claims that models trained in similar ways will develop the same features and circuits, whereas weak universality argues that fundamental principles, such as similar functionalities, are learned across models, but their exact implementations may vary considerably. From this perspective, strong universality would correspond more with "true features", while our paper would align closer to measuring weak universality in terms of similar feature space correlations.
@@
The Importance of Analogous Features. While features across models may not capture the same ground truth feature representations, having analogous feature spaces could suggest that there may be transformations or mappings between the representation spaces of models which reveal their similarity. As such, this has implications for transfer learning and model stitching in which representations, including cross-layer circuits, from one model may be transferred to another (possibly larger) model through a mapping model [Bansal et al., 2021], allowing for feature functionality transfer. These studies may also have implications for furthering studies in neuroscience, as past research has been conducted on measuring the extent of alignment across biological and artificial neural networks [Goldstein et al., 2024], and previous work has also found high-low feature detectors in both vision models [Schubert et al., 2021] and mouse brains [Ding et al., 2023].
@@
This paper seeks to measure the extent of universality in terms of "analogous features in representational space", and unlike Huh et al. [2024], does not claim that models converge towards finding universal features. This follows previous papers which also measured the extent to which different neural networks learn the same representation [Wang et al., 2018a]. We provide evidence that shows there are non-negligible signals of universality captured by subsets of SAE feature spaces across models, but also claim that many feature space subsets in SAEs are dissimilar in terms of their representations and relations in feature space. Still, since this may be due to the limitations of SAEs in capturing features [Engels et al., 2024], these results do not conclude that in these dissimilar areas, models do not share universal features, and further work must be done in this area to draw stronger conclusions.
@@ Appendix I Illustrated Steps of Comparison Method
Figure 19 demonstrates steps 1 to 3 to carry out our similarity comparison experiments. Since SVCCA and RSA are carried out by using the features (or data samples) on the rows, we take the transpose of the decoder weight matrices, which represent the features on its columns.
Note that while these measures are permutation-invariant in their columns, they are not invariant in regards to their rows, which must still be paired correctly [Klabunde et al., 2023]. Additionally, these measures require both matrices to have the same number of rows. For SAEs with the same number of feature columns, this is not an issue, but for SAEs with different number of columns, we only consider 1-1 mappings.
@@
Appendix L Concept Categories List and Further Semantic Subspace Analysis
The keywords used for each concept group for the semantic subspace experiments are given in Table 4. We generate keywords using GPT-4 [Bubeck et al., 2023] using the prompt "give a list of N single token keywords as a python list that are related to the concept of C", such that N is a number (we use N = 50 or 100) and C is a concept like Emotions. We then manually filter these lists to avoid using keywords which have meanings outside of the concept group (eg. "even" does not largely mean "divisible by two" because it is often used to mean "even if…"). We also aim to select case-insensitive keywords which are single-tokens with no spaces.  [^4] However, not all the tokens in the Table 4 are single token.
When searching for keyword matches from each feature’s list of top 5 highest activating tokens, we use keyword matches that avoid features with dataset samples which use the keywords in compound words. For instance, when searching for features that activate on the token "king" from the People/Roles concept, we avoid using features that activate on "king" when "king" is a part of the word "seeking", because "seeking" is unrelated to People/Roles.
We perform interpretability experiments to check, of the features kept, which keywords are activated on. We find that many feature pairs activate on the same keyword and are monosemantic. Most keywords in each concept group are not kept. Additionally, for the layer pairs we checked on, after filtering there were only a few keywords that multiplied features fired on. This shows that the high similarity is not because the same keyword is over-represented in a space.
For instance, for Layer 3 in Pythia-70m vs Layer 5 in Pythia-160m, and for the concept category of Emotions, we find which keywords from the category appear in the top 5 activating tokens of each of the Pythia-70m features from the 24 feature mappings described in Table 1. We only count a feature if it appears once in the top 5 of a feature’s top activating tokens; if three out of five of a feature’s top activating tokens are "smile", then smile is only counted once. The results for Models A and B are given in Table 3. Not only are their counts similar, indicating similar features, but there is not a great over-representation of features. There are a total of 24 features, and a total of 25 keywords (as some features can activate on multiple keywords in their top 5). We note that even if a feature fires for a keyword, that does not mean that that feature’s purpose is only to "recognize that keyword", as dataset examples as just one indicator of what a feature’s purpose is (which still can have multiple, hard-to-interpret purposes).